OpenStudy (nubeer):

integral (cosx/((cosx)^3)+(sinx)^0.5))

OpenStudy (anonymous):

is it\[\sin x^{\frac{1}{2}} \]OR\[(sinx)^{\frac{1}{2}}\] because i tried to integrate the first one worked with me, the other one didnt

OpenStudy (anonymous):

whats that E :O

OpenStudy (nubeer):

(sinx)^1/2

OpenStudy (nubeer):

\[\sqrt{sinx}\]

OpenStudy (mr.math):

Integral of \[\sqrt{\sin x}\] can't be expressed in terms of elementary functions.

OpenStudy (nubeer):

so what is suppose to be done with this question?

OpenStudy (mr.math):

The first term is very easy. It's nothing but \[\int\limits_{}^{}\sec^2x=\tan x.\] As for the second term you can write that it can't be written in terms of the functions you're familiar with. Its integration is just another integration. (Read the definition of the Elliptic Integral in the link above.)

OpenStudy (nikvist):

\[\int\frac{\cos{x}}{\cos^3{x}+\sqrt{sin{x}}}dx\quad\mbox{is that problem?}\]

OpenStudy (anonymous):

oh that makes sense ^